For the overall hypothetical reaction A + 5B → 4C, the rate of appearance of C given by Δ[C]/Δt is the same as

Select one:

a. Δ[A]/Δt

b. -(5/4)(Δ[B]/Δt)

c. -(4/5)(Δ[B]/Δt)

d. none of the above
Answer #1

A + 5B → 4C
The rate is represented as given below [in terms of product and reactants]
-1 *d[A]/dt = - 1/5*d[B]/dt = + 1/4d[C]/dt

- 1/5*d[B]/dt = + 1/4d[C]/dt rearranging this we get

Δ[C]/Δt = -(4/5)(Δ[B]/Δt)

Answer
option C
*********